# Task Graph offloading via Deep Reinforcement Learning in Mobile Edge Computing

### v1

@@ 3 Problem Formulation
If task $v_{ni}$ is scheduled to ECD $m$ under the strategy $\textbf{x}_{ni}$, we can define the execution time $E(x_{ni}^{m})$ of $v_{ni}$ on ECD $m$ as follows.
@@
Except for the aforementioned cases, $v_{ni}$ and $v_{nj}$ are scheduled to ECDs $m$ and $m^{\prime}$ ($m\neq m^{\prime}$) which can communicate directly with each other. We have
@@
Constraint (1) denotes that the decision variable. Constraint (2) claims that the dummy tasks must be located in MU $n$, i.e., ECD 00. Constraint (3) expresses that task $v_{ni}$ can only be scheduled to one ECD.
We can reduce Problem $\mathbf{P}$ to the multiprocessor scheduling problem which is a well-know NP-hard problem garey1979computers . Moreover, it is more difficult to obtain the optimal solution as a result of the dynamic and variable nature of the network environment. Therefore, obtaining the optimal solution to Problem $\mathbf{P}$ will consume high time complexity. For any MU, it needs an efficient strategy with polynomial time rather than the optimal solution, due to the low latency requirement. In this work, we change the optimal control of Problem $\mathbf{P}$ into a Markov Decision Process (MDP) and introduce deep reinforcement learning to solve it.
@@ 4.1 Preparation for MDP
To model a MDP, we must abstract the scheduling of the task graph into a sequence of discrete time steps. At each time step, some representation of the environment can be received by the proposed scheduling mechanism. In the following we present the modeling of discrete time steps for the task graph scheduling.
@@
where $\hat{\delta}$ and $\hat{B}$ are the maximum processing capability in all ECDs and the maximum transmission rate among all ECDs, respectively. The calculation of $\hat{\delta}$ and $\hat{B}$ does not take into account the devices of mobile uses. As the MEC system can serve multiple MUs, each arrived application needs to be calculated by the agent based on Eq. (16). Therefore, we construct a list for the application arriving at different ECDs, in which the application tasks are sorted by the descending order of the task priority. Obviously, the number of lists is equal to the number of applications to be processed concurrently by the system. We denote the set of the lists as =$\mathcal{H}=\{\xi_{n}\}$ ($\forall n\in\mathcal{M}$). Additionally, as the execution of the task graph is constrained by the dependencies among tasks, the scheduling plan of a task can be made after its parents have been made. We call such a task the ready task.
When the agent receives the arrival event or the completion event, it will pick out the ready tasks from $\xi_{n}$ and arrange them into the queue $Q^{r}$ in an ascending order of their priorities. Subsequently, the agent will fetch the first ready task from $Q^{r}$, in turn, and make the scheduling plan for it. Built upon such a mechanism, we think of the time when the agent gets each task out of $Q^{r}$ at a time step. The agent will interact with the environment to plan for task scheduling via a reinforcement learning algorithm at each time step.
@@ 4.2 MDP Mode of Agent
MDP mode requires that the agent can receive reward signals by observing the state of the environment. The decision controller can periodically collect real-time status information about the system so that the agent can receive reward signals from the state of the environment. The reward signals are used to evaluate the quality of actions available to agents. Then, the agent selects an action based on the evaluation. Afterward, at the next time step the environment changes into a new state. The agent can receive new reward signals according to the feedback of the environment. The agent chooses actions with the goal of maximizing the sum of the expected reward obtained at each time step. Next, we identify the state space, the action space and the reward function in our MDP as follows.
1) State Space: In a MDP the environment can be abstracted into a state space, which will change from one state to another after an action is selected by the agent. We define the state of MDP corresponding to the MEC system at time step $\tau$ as a vector $\textbf{s}_{\tau}=(\hat{\hat{B}},B_{n}^{m},\hat{\hat{\delta}},w^{r},w^{m},w^{n})$, which contains several parameters about computing delay and transmission delay. Specifically, $\hat{\hat{B}}$ denotes the sum of transmission rates between different ECDs in the network, i.e., $\hat{\hat{B}}=\sum\nolimits_{m=1}^{M}\sum\nolimits_{m^{\prime}=1}^{M}B_{m,m^{\prime}}$; $\hat{\hat{\delta}}$ expresses the sum of processing capacities of all ECDs in the network, i.e., $\hat{\hat{\delta}}=\sum\nolimits_{m=1}^{M}\delta_{m}$; $w^{r}$ indicates the total workloads of the tasks in the ready queue $Q^{r}$; $w^{m}$ is the total workloads of the tasks in computing queues of all ECDs; $w^{n}$ is the total workloads of the tasks in the computing queue of MU $n$. Note that, all parameters in $\textbf{s}_{\tau}$ can be observed from the real-time status information.
2) Action Space: The MDP uses the action to describe the behavior of an agent. Specifically, the action is the decision made by the agent after watching the state at a certain time step $\tau$. We use $\textbf{a}_{\tau}$ to denote the action space available to the agent when making scheduling decisions for each task. As each task except for the two dummy tasks can be scheduled to one of the ECDs in the network, $\textbf{a}_{\tau}=(a^{0},a^{1},\ldots,a^{M})$ expresses the strategy space to schedule task $v_{ni}$. The item $a^{m}$ $(\forall m\in\{0,1,\ldots,M\})$ is the binary variable so that $\textbf{a}_{\tau}$ corresponds to the strategy $\textbf{x}_{ni}$ defined in Section LABEL:systemModel.
@@ 4.3 Deep Reinforcement Learning
DQN algorithm can improve the search speed of the traditional Q-Learning by introducing the deep neural network, called the prediction network. The prediction network is responsible for estimating the Q-value at state $\textbf{s}_{\tau}$. The algorithm selects an action to make a scheduling decision. To improve the algorithm’s stability and diversity, DQN sets the experience pool and adds the target network. The structure of the target network is consistent with that of the prediction network. The experience pool is the memory space that stores the records of the transition samples captured by the agent and the environment at each time step. Each record consists of four items ($\textbf{s}_{\tau},\textbf{a}_{\tau},\textbf{r}_{\tau+1},\textbf{s}_{\tau+1}$), i.e., the state, the selected action, the received reward and the subsequent state. DQN extracts a batch of samples randomly from the experience pool to improve the data correlation and non-static distribution. Specifically, a batch of elements $\textbf{s}_{\tau}$ in the transition records is used as the input to the prediction network. Meanwhile, a batch of elements $\textbf{s}_{\tau+1}$ in the transition records is transmitted to the target network. During the training process, the parameters of the prediction network are copied to the target network after a certain number of round iterations. In DQN, the Q-value of the training process generated by the prediction network is called the current Q-value, and that of the training process generated by the target network is called the target Q-value. DQN maintains the difference between the parameters of the two neural networks to calculate the loss function. Whereafter, the parameters of the prediction network are inversely updated by leveraging the stochastic gradient descent (SGD). We define the loss function as follows.
@@ 5.2 Resource Allocation Algorithm for Task Graph
The proposed algorithm consists of the scheduling algorithm for the task graph, named SATA, and the deep reinforcement learning algorithm, called DRL. SATA algorithm undertakes the preparation for MDP. It takes charge of fetching out all the ready tasks from $\mathcal{H}$ and pushing them into queue $Q^{r}$ to invoke DRL on receipt of the aforementioned two events.
Algorithm 1 shows the processing of the task graph scheduling. This algorithm is launched on receipt of the arrival and completion events. In Lines 2-7 the tasks of an arrived application $n$ are sorted in list $\xi_{n}$ by a descending order of $F_{ni}^{lct}$, meanwhile, and list $\xi_{n}$ is appended to $\mathcal{H}$, when the arrival event triggers. This algorithm’s Lines 8-19 will be launched when either the arrival event or the completion event triggers. Specifically, Lines 8-11 express that all ready tasks in the set $\mathcal{H}$ are mapped into $Q^{r}$ by an ascending order of the latest completion time. In Lines 12-19 SATA generates the sequence of discrete time steps for the task graph scheduling by iteratively fetching out a ready task $v_{ni}$ from $Q^{r}$. SATA obtains the current state $\textbf{s}_{\tau}$ from the environment at each time step $\tau$. Notice that, if the fetched task is the first of all application tasks, the reward is a null value. Otherwise, SATA will calculate the current reward $\textbf{r}_{\tau}$ based on Eqs. (17),(18),(19) according to the environment information. In Line 17 SATA passes $\textbf{s}_{\tau}$ and $\textbf{r}_{\tau}$ to the deep reinforcement learning algorithm, and subsequently receives the action space $\textbf{a}_{\tau}$ from it. Thus, the task $v_{ni}$ can be scheduled to the target ECD in the light of $\textbf{a}_{\tau}$. The SATA algorithm will end when no new applications arrive or no tasks are completed.
@@
6:        Rank all tasks in list $\xi_{n}$ by a descending order of $F_{ni}^{lct}$;
@@
For the DRL algorithm, it is to implement deep reinforcement learning. Algorithm 2 shows the processing that SATA makes the scheduling decision based on deep reinforcement learning. This algorithm first constructs two neural networks with stochastic parameters $\theta$ and $\theta^{-}$, respectively. Then, it will receive state $\mathcal{S}$ and reward $\mathcal{R}$ from the SATA algorithm at each time step. The prediction network outputs the Q-value by inputting state $\mathcal{S}$. Based on the $\varepsilon$-greedy exploration, action $\mathcal{A}$ is chosen out. The above processes are described in lines 4-7 of this algorithm. Notice that, if the current time step is processing the first task of all applications to be scheduled, i.e., $\tau=1$, the algorithm will skip the storage of the experience pool and the computation of the target network and prepare the state and action for storing the experience transition in the next time step. So, Lines 8-16 are skipped and Line 17 is run directly when $\tau=1$. When $\tau>1$, the transition consisting of the last state $\textbf{s}_{\tau}$ and action $\textbf{a}_{\tau}$ as well as the current $\textbf{r}_{\tau+1}$ and $\textbf{s}_{\tau+1}$ is stored into the experience pool $\mathbb{D}$, as shown in Line 10. Then, the DRL algorithm will sample a batch $\mathcal{B}$ of transitions from $\mathbb{D}$ when the number of time steps is more than the predefined batch size, as shown in Lines 11-12. For each transition in $\mathcal{B}$, the DRL algorithm needs to calculate the loss function based on the output from the prediction network and target network, respectively, and finally performs SGD to update the prediction network. These pseudo-codes are shown in Lines 13-16.
@@
4:     Receive state $\mathcal{S}$ and reward $\mathcal{R}$ from SATA algorithm at step $\tau$;
@@
7:     Pass $\mathcal{A}$ to SATA algorithm;
@@
17:     $\textbf{s}_{\tau}\leftarrow S$; $\textbf{a}_{\tau}\leftarrow\mathcal{A}$;
@@ 5.3 Intelligent Agent Training and Working
6:        Rank all tasks in list $\xi_{n}$ by a descending order of $F_{ni}^{lct}$;
@@ 6 Performance evaluation
In this section we evaluate SATA-DRL algorithm in the simulator built by combining ElasticSimcai2017elasticsim and EdgeCloudSimsonmez2018edgecloudsim . Moreover, we compare SATA-DRL with several competing algorithms for the evaluation of task graph scheduling in the MEC system.
@@ 6.1 Simulation Setup
In our experiments, we let the processing capabilities of four ECDs transition from one state to another according to this probability matrix after a task is completed on any computing device. Besides, the transmission rate between ECDs is set to 440 Mbps, and that between each MU and ECD $m_{n}$, the ECD covering the MU, is set to $10^{3}$ Mbps. These two transmission rates are similar to the literature sundar2018offloading . In addition, $\hat{\delta}$ and $\hat{B}$ are set to 6000 MIPS and $10^{3}$ Mbps, respectively.
We utilize one type of workflow datasets juve2013characterizing , i.e., Montage with 25 nodes, to simulate the task graph of each UT’s application. The workload of a task in an application is set to 500 if the original value of the element $runtime$ recorded in the DAX file is greater than 500. Otherwise, that of a task is set to 100 if the original $runtime$ is less than 100. To set the size of transfer data, we need to calculate the average transmission rate $\overline{B}$ as $(440\times 6+10^{3})/7=520$ (Mbps) in the MEC system. Then, based on the element $size$ recorded in the DAX file, we can calculate a base communication time, named $bc$. We set $bc$ to $size/\overline{B}$, respectively, if $bc=size/\overline{B}$ does not exceed the range of $[10^{-3},10^{-2}]$. Otherwise, let $bc$ equal to $10^{-3}$ and $10^{-2}$, respectively. Based on this, we set the size of the data transfer between tasks in an application to $bc\cdot\overline{B}$. The size of the data transfer associated with the dummy tasks is set to $bc\cdot\overline{B}$.

### v4

@@ 4 Problem Formulation
For the other real tasks $v_{ni}$, i.e., $i\neq\{0,I\}$, they need to be offloaded to the MEC system rather than executed locally. Thus, we have
$x_{ni}^{0}=0,\quad i\neq\{0,I\}.$ (3)
@@
If task $v_{ni}$ is scheduled to ECD $m$ under the strategy $\textbf{x}_{ni}$, we can define the execution time $E(\textbf{x}_{ni})$ of $v_{ni}$ on ECD $m$ as follows.
@@
Except for the aforementioned cases, $v_{ni}$ and $v_{nj}$ are respectively scheduled to real ECDs $m$ and $m^{\prime}$ ($m\neq m^{\prime}$) which can communicate directly with each other. We have
@@
Constraint (1) denotes that the decision variable. Constraint (2) claims that the dummy tasks must be located in MU $n$, i.e., ECD 00. Constraint (3) indicates that all real tasks are offloaded to the MEC system. Constraint (4) expresses that task $v_{ni}$ can only be scheduled to one ECD.
We can reduce Problem $\mathbf{P}$ to the multiprocessor scheduling problem which is a well-know NP-hard problem lewis1983michael . Moreover, it is more difficult to obtain the optimal solution as a result of the dynamic and variable nature of the network environment. Therefore, obtaining the optimal solution to Problem $\mathbf{P}$ will consume high time complexity. For any MU, it needs an efficient strategy with polynomial time rather than the optimal solution, due to the low latency requirement. The heuristic or approximation methods can find a feasible solution to task offloading rapidly, but they rely heavily on expert knowledge or accurate analytical models wang2021dependent . It is also more difficult to formulate an accurate analytical model in a dynamic environment, which is time-consuming due to the increasing complexity of applications wang2021dependent . DRL can make sequential decisions by self-learning in a dynamic environment, without the need for expert knowledge. In this work, we change the optimal control of Problem $\mathbf{P}$ into a Markov Decision Process (MDP) and introduce deep reinforcement learning to solve it.
@@ 5.1 Preparation for MDP
To model a MDP, we must abstract the scheduling of the task graph into a sequence of discrete time steps. At each time step, the representation of the environment can be received by the proposed scheduling mechanism. In the following we present the modeling of discrete time steps for the task graph scheduling.
@@
where $\hat{\delta}$ and $\hat{B}$ are the maximum processing capability in all ECDs and the maximum transmission rate among all ECDs, respectively. The calculation of $\hat{\delta}$ and $\hat{B}$ does not take into account the devices of mobile uses. As the MEC system can serve multiple MUs, each arrived application needs to be calculated by the agent based on Eq. (17). Therefore, we construct a list $\xi_{n}$ for each application $n$ arriving at different ECDs, in which the application tasks are sorted by the ascending order of the task priority. Obviously, the number of lists is equal to the number of applications to be processed concurrently by the system. We denote the set of the lists as =$\mathcal{H}=\{\xi_{n}\}$ ($\forall n\in\mathcal{N}$).
As the execution of the task graph is constrained by the dependencies among tasks, the scheduling plan of a task can be made after its parents have been made. We call such a task the ready task. When the agent receives the arrival event or the completion event, it will pick out the ready tasks from $\xi_{n}$ and arrange them into the queue $Q^{r}$ in an ascending order of their priorities. Subsequently, the agent will fetch the first ready task from $Q^{r}$, in turn, and make the scheduling plan for it. Built upon such a mechanism, we think of the time when the agent gets each task out of $Q^{r}$ at a time step. The agent will interact with the environment to plan for task scheduling via a reinforcement learning algorithm at each time step.
@@ 5.2 MDP Mode of Agent
The MDP mode requires that the agent can receive reward signals by observing the state of the environment. The decision controller can periodically collect real-time status information about the system so that the agent can receive reward signals from the state of the environment. The reward signals are used to evaluate the quality of actions available to the agent. Then, the agent selects an action based on the evaluation. Afterward, at the next time step the environment changes into a new state. The agent can receive new reward signals according to the feedback of the environment. The agent chooses actions with the goal of maximizing the sum of the expected reward obtained at each time step. Next, we represent the state space, the action space and the reward function in our MDP as follows.
1) State Space: In a MDP the environment can be abstracted into a state space, which will change from one state to another after an action is selected by the agent. We define the state of MDP corresponding to the MEC system at time step $\tau$ as a vector $\textbf{s}_{\tau}=(\hat{\hat{B}},B_{n}^{m},\hat{\hat{\delta}},w^{r},w^{m})$, which contains five parameters about computing delay and transmission delay. Specifically, $\hat{\hat{B}}$ denotes the sum of transmission rates between different ECDs in the network, i.e., $\hat{\hat{B}}=\sum\nolimits_{m=1}^{M}\sum\nolimits_{m^{\prime}=1}^{M}B_{m,m^{\prime}}$; $\hat{\hat{\delta}}$ expresses the sum of processing capacities of all ECDs in the network, i.e., $\hat{\hat{\delta}}=\sum\nolimits_{m=1}^{M}\delta_{m}$; $w^{r}$ indicates the total workloads of the tasks in the ready queue $Q^{r}$; $w^{m}$ is the total workloads of the tasks in computing queues of all ECDs. Note that, all parameters in $\textbf{s}_{\tau}$ can be observed by the agent from the real-time status information.
2) Action Space: The MDP uses the action to describe the behavior of an agent. Specifically, the action is the decision made by the agent after watching the state at a certain time step $\tau$. We use $\textbf{a}_{\tau}$ to denote the action space available to the agent when making scheduling decisions for each task. As each task except for the two dummy tasks can be scheduled to one of the ECDs in the network, $\textbf{a}_{\tau}=(a^{0},a^{1},\ldots,a^{M})$ expresses the strategy space to schedule task $v_{ni}$. The item $a^{m}$ $(\forall m\in\{0,1,\ldots,M\})$ is the binary variable so that $\textbf{a}_{\tau}$ corresponds to the strategy $\textbf{x}_{ni}$ defined in Section 4.
@@ 5.3 Deep Reinforcement Learning
DQN algorithm can improve the search speed of the traditional Q-Learning by introducing the deep neural network, called the prediction network. The prediction network is responsible for estimating the Q-value at state $\textbf{s}_{\tau}$. The algorithm selects an action to make a scheduling decision. To improve the algorithm’s stability and diversity, DQN sets the experience pool and adds the target network. The structure of the target network is consistent with that of the prediction network. The experience pool is the memory space that stores the records of the transition samples captured by the agent from the environment at each time step. Each record consists of four items ($\textbf{s}_{\tau},\textbf{a}_{\tau},\textbf{r}_{\tau+1},\textbf{s}_{\tau+1}$), i.e., the state, the selected action, the received reward and the subsequent state. DQN extracts a batch of samples randomly from the experience pool to improve the data correlation and non-static distribution. Specifically, a batch of elements $\textbf{s}_{\tau}$ in the transition records is used as the input to the prediction network. Meanwhile, a batch of elements $\textbf{s}_{\tau+1}$ in the transition records is transmitted to the target network. During the training process, the parameters of the prediction network are copied to the target network after a certain number of round iterations. In DQN, the Q-value of the training process generated by the prediction network is called the current Q-value, and that of the training process generated by the target network is called the target Q-value. DQN maintains the difference between the parameters of the two neural networks to calculate the loss function. Whereafter, the parameters of the prediction network are inversely updated by leveraging the stochastic gradient descent (SGD). We define the loss function as follows.
@@ 6.2 Resource Allocation Algorithm for Task Graph
We design a resource scheduling algorithm, named SATA, for the task graph based on an event-driven mechanism. It can transform the information received from the environment into the state and reward in the MDP model. Then, the state and reward will be passed to the deep reinforcement learning algorithm, called DRL. SATA takes charge of fetching out all the ready tasks from $\mathcal{H}$ and pushing them into queue $Q^{r}$ to invoke DRL.
Algorithm 1 shows the processing of the task graph scheduling. This algorithm is launched on receipt of the arrival and completion events. In Lines 2-7 the tasks of an arrived application $n$ are sorted in list $\xi_{n}$ by an ascending order of $F_{ni}^{lct}$, meanwhile, and list $\xi_{n}$ is appended to $\mathcal{H}$ when the arrival event triggers. This algorithm’s Lines 8-19 will be launched when either the arrival event or the completion event triggers. Specifically, Lines 8-11 express that all ready tasks in the set $\mathcal{H}$ are mapped into $Q^{r}$ by an ascending order of the latest completion time. In Lines 12-19 SATA generates the sequence of discrete time steps for the task graph scheduling by iteratively fetching out a ready task $v_{ni}$ from $Q^{r}$. SATA obtains the current state $\textbf{s}_{\tau}$ from the environment at each time step $\tau$. Notice that, if the fetched task is the first of all application tasks, the reward is a null value. Otherwise, SATA will calculate the current reward $\textbf{r}_{\tau}$ based on Eqs. (18),(19),(20) according to the environment information. In Line 17 SATA passes $\textbf{s}_{\tau}$ and $\textbf{r}_{\tau}$ to the deep reinforcement learning algorithm, and subsequently receives the action space $\textbf{a}_{\tau}$ from it. Thus, the task $v_{ni}$ can be scheduled to the target ECD in the light of $\textbf{a}_{\tau}$. The SATA algorithm will end when no new applications arrive or no tasks are completed.
@@
6:        Rank all tasks in list $\xi_{n}$ by an ascending order of $F_{ni}^{lct}$;
@@
For the DRL algorithm, it is to implement deep reinforcement learning. Algorithm 2 shows the processing that SATA makes the scheduling decision based on deep reinforcement learning. This algorithm first constructs two neural networks with stochastic parameters $\theta$ and $\theta^{-}$, respectively. Then, it will receive state $\mathcal{S}$ and reward $\mathcal{R}$ from the SATA algorithm at each time step. The prediction network outputs the Q-value by inputting state $\mathcal{S}$. Based on the $\varepsilon$-greedy exploration, action $\mathcal{A}$ is chosen out. The above processes are described in lines 4-7 of this algorithm. Notice that, if the current time step is processing the first task of all applications to be scheduled, i.e., $\tau=1$, the algorithm will skip the storage of the experience pool and the computation of the target network, and it will prepare the state and action for storing the experience transition in the next time step. So, Lines 8-16 are skipped and Line 17 is run directly when $\tau=1$. When $\tau>1$, the transition consisting of the last state $\textbf{s}_{\tau}$ and action $\textbf{a}_{\tau}$ as well as the current $\textbf{r}_{\tau+1}$ and $\textbf{s}_{\tau+1}$ is stored into the experience pool $\mathbb{D}$, as shown in Line 10. Then, the DRL algorithm will sample a batch $\mathcal{B}$ of transitions from $\mathbb{D}$ when the number of time steps is more than the predefined batch size, as shown in Lines 11-12. For each transition in $\mathcal{B}$, the DRL algorithm needs to calculate the loss function based on the output from the prediction network and target network, respectively, and it finally performs SGD to update the prediction network. These pseudo-codes are shown in Lines 13-16.
@@
4:     Receive state $\mathcal{S}$ and reward $\mathcal{R}$ from the SATA algorithm at step $\tau$;
@@
7:     Pass $\mathcal{A}$ to the SATA algorithm;
@@
17:     $\textbf{s}_{\tau}\leftarrow\mathcal{S}$; $\textbf{a}_{\tau}\leftarrow\mathcal{A}$;
@@ 6.3 Intelligent Agent Training and Working
6:        Rank all tasks in list $\xi_{n}$ by an ascending order of $F_{ni}^{lct}$;
@@ 7 Performance evaluation
To evaluate SATA-DRL, we extend CloudSim calheiros2011cloudsim according to two simulators, i.e., EdgeCloudSim sonmez2018edgecloudsim and ElasticSim cai2017elasticsim . Specifically, we tailor CloudSim to the problem of dependent task offloading by combining two approaches as follows. The one is the simulation of edge service devices supported by the EdgeCloudSim simulator. The other is the generation of workflows supported by the ElasticSim simulator. Moreover, we compare SATA-DRL with several competing algorithms for the evaluation of task graph scheduling in the MEC system.
@@ 7.1 Simulation Setup
In our experiments, we let the processing capabilities of four ECDs transition from one state to another according to this probability matrix after a task is completed on any computing device. Besides, the transmission rate between ECDs is set to 440 Mbps, and that between each MU and ECD $m_{n}$, i.e., the ECD covering the MU, is set to $10^{3}$ Mbps. These two transmission rates are similar to the literature sundar2018offloading . In addition, $\hat{\delta}$ and $\hat{B}$ are set to 6000 MIPS and $10^{3}$ Mbps, respectively.
The scientific workflow datasets juve2013characterizing contain the workflow characterizations corresponding to task graph applications. It can provide a community resource to evaluate workflow algorithms. As Montage workflow has been used extensively to evaluate workflow algorithms and systems, the literature zhang2018resource utilized the Montage workflow to evaluate its task graph offloading algorithm. Similar to zhang2018resource , We choose the Montage type of workflow datasets, which contains 25 nodes, to simulate the task graph of each UT’s application. The workload of a task in an application is set to 500 if the original value of the element $runtime$ recorded in the DAX file is greater than 500. Otherwise, that of a task is set to 100 if the original $runtime$ is less than 100. To set the size of transfer data, we need to calculate the average transmission rate $\overline{B}$ as $(440\times 6+10^{3})/7=520$ (Mbps) in the MEC system. Then, based on the element $size$ recorded in the DAX file, we can calculate a base communication time, named $bc$. We set $bc$ to $size/\overline{B}$, if $bc=size/\overline{B}$ does not exceed the range of $[10^{-3},10^{-2}]$. Otherwise, let $bc$ equal to $10^{-3}$ and $10^{-2}$, respectively. Based on this, we set the size of the data transfer between tasks in an application to $bc\cdot\overline{B}$. The size of the data transfer associated with the dummy tasks is set to $bc\cdot\overline{B}$.